# A Module-theoretic Introduction to Abstract Elementary Classes

### v1

@@ 3.2.1. Stability in Model Theory
    Syntactic stability: There is some $\lambda$ so that, for any $M\vDash T$ of size $\lambda$, $|S(M)|=\lambda$.
@@ 3.2.2. Galois stability in modules
The most basic question about stability in AECs of modules is wether the analogue of the Baur-Fisher result holds for Galois stability, first asked by Mazari-Armida [MA21a, Question 2.12].
@@
As listed, there are two common methods of proof. First, if one camn prove that Galois types are syntactic for some fragment of first-order logic (typically either quantifier-free of p.p. formulas), then the number of Galois types is bounded above by the number of first-order syntactic types, so Galois stability follows from the classic Baur-Fischer result for first-order. Second, some of the properties of a stable independence relation can be used to prove Galois stability using a counting argument.
Note that the first-order Baur-Fisher result deals with the number of syntactic types in a complete first-order theory of modules. While some of the examples are refinements of these elementary classes (Fact 3.14.(2) for instance), others are not (Fact 3.14.(1) for instance). There is some subtle interplay of monotonicity when it comes to Galois types, but it is not the direction that helps answer the question. For instance, given two classes $\mathbb{K}^{0},\mathbb{K}^{1}$, modules $M\subset_{pr}N_{0},N_{1}$, and elements $a_{\ell}\in N_{\ell}$, we have
@@ 3.3.1. Limit models in Abstract Elementary Classes
    $M\in\mathbb{K}_{\lambda}$ is a $(\lambda,\alpha)$-limit model iff there is a continuous $\prec_{\mathbb{K}}$-increasing chain $\{M_{\alpha}\in\mathbb{K}_{\lambda}:\alpha<\lambda^{+}\}$ such that
@@
    If $\mathbb{K}_{\lambda}$ has amalgamation and is Galois stable in $\lambda$, then for every $M\in\mathbb{K}_{\lambda}$, there is universal $N\in\mathbb{K}_{\lambda}$ so $M\prec N$.
@@
    Suppose that for every $M\in\mathbb{K}_{\lambda}$, there is universal $N\in\mathbb{K}_{\lambda}$ so $M\prec N$. Then $\mathbb{K}_{\lambda}$ has amalgamation and is Galois stable in $\lambda$.
@@
The answer to this riddle is the union of saturated models. Saturated models (see [Bal09, Definition 8.13], especially $M$ is saturated iff it is $\|M\|$-saturated) are well-studied in stable theories. In particular , for a stable theory $T$, there is an important cardinal $\kappa(T)<|T|^{+}$ with several different meanings, including measuring the local character of the resulting independence relation (so smaller $\kappa(T)$ indicate a stronger independence relation).
Another way of understanding $\kappa(T)$ is to look at in increasing chain $\{M_{i}:i<\alpha\}$ of $\lambda$-saturated models and ask if the union $\bigcup_{i<\alpha}M_{i}$ is also $\lambda$-saturated. A first approximation would be to compare $\lambda$ and $\alpha$: if $\lambda<\text{cf }\alpha$, then any $<\lambda$-sized set of parameters from the union appears already in some $M_{i}$; thus, any type over it is realized in $M_{i}$ and the union is $\lambda$-saturated. However, this first approximation could never show that a short union of highly saturated models is highly saturated.
@@
Now we return to Abstract Elementary Classes and the uniqueness of limit models. We infer from first order that the union of saturated models being saturated is an important concept to explore in nonelementary classification theory (and this has been shown true in works such as [BV17a]). However, the local analysis of Abstract Elementary Classes does not allow us to use saturated models! We work with Galois types over models and restricting our focus to $\mathbb{K}_{\lambda}$ means there are no smaller models to look at types over.
@@
Suppose $\mathbb{K}$ is an Abstract Elementary Class and $\lambda>\text{LS}(\mathbb{K})$ so $\mathbb{K}_{\lambda}$ has amalgamation and is Galois stable in $\lambda$.
@@
    $\mathbb{K}_{\lambda}$-satisfies the uniqueness of limit models iff the union of Galois saturated models in $\mathbb{K}_{\lambda}$ of length $<\lambda^{+}$ is Galois saturated.
@@ 3.3.2. Limit models in classes of modules
Prior to Mazari-Armida’s work [MA20], the study of limit models and their uniqueness had been largely theoretical: it was explored in abstract set ups, but no examples had been explored. [MA20] examined the class of torsion-free abelian groups and showed that a $(\lambda,\alpha)$-limit model with cf $\text{cf }\alpha\geq\omega_{1}$ must be algebraically compact [MA20, Lemma 4.10] owing to the uncountable cofinality of $\alpha$ and the countable condition on algebraic compactness, which then allows one to deduce that all such limits must have the same structure [MA20, Lemma 4.14]. On the other hand, $(\lambda,\omega)$-limit models are isomorphic to the other models (and thus not algebraically compact). The argument for this relies heavily on model theoretic work: if all limit models were isomorphic, then this gives a form of superstability, namely [MA20, Fact 4.19] combines work of Grossberg and Vasey [Vas16, GV17] shows that the nice AEC conditions that $\mathbb{K}^{tf}$ satisfies plus uniqueness of limit models would imply Galois stability on a tail of cardinals; however, $\mathbb{K}^{tf}$ has long been known to be only strictly stable (in particular, stable in all $\lambda=\lambda^{\aleph_{0}}$, see [BET07]).
Similarly, $\mathbb{K}^{tor}$ is shown [MA23b, Lemma 5.7] to be strictly stable by showing $(\lambda,\omega)$-limir and $(\lambda,\omega_{1})$-limit models are not isomorphic by giving explicit descriptions of them.
Further work on limit models picked up on a familiar theme in commutative algebra: determining properties of the ring $R$ by understanding properties of $R$-modules.
@@
The proofs of these different results are different. One neat commonality is that the proof of uniqueness of limit models often goes through a Schroeder-Bernstein style argument: for instance, in 1) and (3), one shows that any two limit models of the are injective and that they are embeddable into each other because they are universal, then uses a result of Bumby (see [MA21b, Fact 2.26]) to conclude they are isomorphic.

### v2

@@ 3.2.1. Stability in Model Theory
    Syntactic stability: There is some $\lambda$ such that, for any $M\vDash T$ of size $\lambda$, $|S(M)|=\lambda$, where $S(M)$ denotes the set of syntactic types over $M$.
@@ 3.2.2. Galois stability in modules
The most basic question about stability in AECs of modules is whether the analogue of the Baur-Fisher result holds for Galois stability, first asked by Mazari-Armida [MA21a, Question 2.12].
@@
As listed, there are two common methods of proof. First, if one can prove that Galois types are syntactic for some fragment of first-order logic (typically either quantifier-free of p.p. formulas), then the number of Galois types is bounded above by the number of first-order syntactic types, so Galois stability follows from the classic Baur-Fischer result for first-order. Second, some of the properties of a stable independence relation can be used to prove Galois stability using a counting argument.
Note that the first-order Baur-Fisher result deals with the number of syntactic types in a complete first-order theory of modules. While some of the examples are refinements of these elementary classes (Fact 3.14.(2) for instance), others are not (Fact 3.14.(1) for instance). There is some subtle interplay of monotonicity when it comes to Galois types, but it is not the direction that helps answer the question. For instance, given two classes $\mathbb{K}^{0}\subset\mathbb{K}^{1}$ and modules $M\subset_{pr}N_{0},N_{1}$ in $\mathbb{K}^{0}$, and elements $a_{\ell}\in N_{\ell}$, we have
@@ 3.3.1. Limit models in Abstract Elementary Classes
    $M\in\mathbb{K}_{\lambda}$ is a $(\lambda,\alpha)$-limit model iff there is a continuous $\prec_{\mathbb{K}}$-increasing chain $\{M_{i}\in\mathbb{K}_{\lambda}:i<\alpha\}$ such that
@@
[…]
@@
[…]
@@
The answer to this riddle is the study of when the union of saturated models is saturated. Saturated models (see [Bal09, Definition 8.13], especially $M$ is saturated iff it is $\|M\|$-saturated) are well-studied in stable theories. In particular, for a stable theory first-order $T$, there is an important cardinal $\kappa(T)<|T|^{+}$ with several different meanings, including measuring the local character of the resulting independence relation (so smaller $\kappa(T)$ indicate a stronger independence relation).
Another way of understanding $\kappa(T)$ is to look at an increasing chain $\{M_{i}:i<\alpha\}$ of $\lambda$-saturated models and ask if the union $\bigcup_{i<\alpha}M_{i}$ is also $\lambda$-saturated. A first approximation would be to compare $\lambda$ and $\alpha$: if $\lambda<\text{cf }\alpha$, then any $<\lambda$-sized set of parameters from the union appears already in some $M_{i}$; thus, any type over it is realized in $M_{i}$ and the union is $\lambda$-saturated. However, this first approximation could never show that a short union of highly saturated models is highly saturated.
@@
Now we return to Abstract Elementary Classes and the uniqueness of limit models. We infer from first order that the union of saturated models being saturated is an important concept to explore in nonelementary classification theory (and this has been shown true in works such as [BV17a]). However, the local approach [^18] The ‘local approach’ and ‘global approach’ are defined and discussed in [BV17b, Section 2.1]. Briefly, the local approach (exemplified by [She01]) starts with assumptions on $\mathbb{K}_{\lambda}$ to prove properties of $\mathbb{K}_{\lambda^{+}}$ and above, while the global approach (exemplified by [She99]) makes assumptions on all of $\mathbb{K}$. to analyzing Abstract Elementary Classes does not allow us to use saturated models! We work with Galois types over models and restricting our focus to $\mathbb{K}_{\lambda}$ means there are no smaller models to look at types over.
@@
[…]
@@
    $\mathbb{K}_{\lambda}$ satisfies the uniqueness of limit models iff the union of Galois saturated models in $\mathbb{K}_{\lambda}$ of length $<\lambda^{+}$ is Galois saturated.
@@
So far, the term ‘superstability’ has not been defined. One reason for this is that the several equivalent characterizations of superstability in elementary classes bifurcate into distinct properties in Abstract Elementary Classes (although [GV17] shows that they reconverge in tame Abstract Elementary Classes). For its most basic definition, superstability is used to mean ‘Galois stable on a tail of cardinals.’. However, one of the other possible definitions of superstability is the uniqueness of limit models, and some sources have begun to use this meaning as its definition.
@@ 3.3.2. Limit models in classes of modules
Prior to Mazari-Armida’s work [MA20], the study of limit models and their uniqueness had been largely theoretical: it was explored in abstract set ups, but no examples had been explored. [MA20] examined the class of torsion-free abelian groups and showed that a $(\lambda,\alpha)$-limit model with cf $\text{cf }\alpha\geq\omega_{1}$ must be algebraically compact [MA20, Lemma 4.10] owing to the uncountable cofinality of $\alpha$ and the countable condition on algebraic compactness, which then allows one to deduce that all such limits must have the same structure [MA20, Lemma 4.14]. On the other hand, $(\lambda,\omega)$-limit models are not isomorphic to the other models (and thus not algebraically compact). The argument for this relies heavily on model theoretic work: if all limit models were isomorphic, then this gives a form of superstability, namely [MA20, Fact 4.19] combines work of Grossberg and Vasey [Vas16, GV17] to show that the nice AEC conditions that $\mathbb{K}^{tf}$ satisfies plus uniqueness of limit models would imply Galois stability on a tail of cardinals; however, $\mathbb{K}^{tf}$ has long been known to be only strictly stable (in particular, stable in exactly the cardinals $\lambda=\lambda^{\aleph_{0}}$, see [BET07]).
Similarly, $\mathbb{K}^{tor}$ is shown [MA23b, Lemma 5.7] to be strictly stable by showing $(\lambda,\omega)$-limit and $(\lambda,\omega_{1})$-limit models are not isomorphic by giving explicit descriptions of them.
Further work on limit models picked up on a familiar theme in ring theory: determining properties of the ring $R$ by understanding properties of $R$-modules.
@@
The proofs of these different results are different. One neat commonality is that the proof of uniqueness of limit models often goes through a Schroeder-Bernstein style argument: for instance, in 1) and (3), one shows that any two limit models of $\mathbb{K}^{\text{R-Mod}}$ are injective and that they are embeddable into each other because they are universal, then uses a result of Bumby (see [MA21b, Fact 2.26]) to conclude they are isomorphic.